A metal cube(faces 20 mm+-0.1mm) fully sinks into a glass full of water. The diameter of the glass is 60 mm+-0,2 mm.Calculate the water rise level and its cumulative error.

1
Expert's answer
2022-01-16T17:56:02-0500
"y=\dfrac{4a^3}{\pi d^2}=\dfrac{4·20^3}{\pi ·60^2}=2.8\text{ mm},\\\space\\
\Delta y=0.1+0.2=0.3 \text{ mm}."